# Dual effect of cholesterol on interfacial water dynamics in lipid membranes: Interplay between membrane packing and hydration

### v1

@@ III.3 Water orientation near membrane interface
We further analyze how the orientation of water molecules at the membrane interface depends on Chol concentration. Specifically, we calculated the angle $\theta$ between the water dipole vector and the bilayer normal. Previous MD simulations have demonstrated that water orientation varies with the nature of the interacting functional groups. [71, 72, 73, 35, 74, 75, 76] While these results typically represent the $\cos\theta$ dependence on distance $z$ from the membrane center, it should be noted that our analysis uses $z^{\prime}$, a coordinate defined relative to the nearest phosphorus atom. This approach accounts for local fluctuations of the membrane interface and, in particular, enables a more accurate characterization of water orientation near the lipid headgroup region.
@@
Figure 5 demonstrates that in Region 3 (bulk region) defined by $z^{\prime}>0.85$ nm, $\rho(z^{\prime},\cos\theta)$ is approximately 0.5 regardless of $\cos\theta$, indicating an isotropic orientation of water molecules. Furthermore, it is demonstrated that both DPPC and PSM exhibit higher values at $-1<\cos\theta<0.5$ in Region 2 defined by $-0.25~\mathrm{nm}<z^{\prime}<0.85~\mathrm{nm}$. This Region 2 likely represents water molecules located near the phosphate groups of the lipid headgroups, as discussed in Sec III.2. The predominantly negative values of $\cos\theta$ suggest that the dipole vectors of these water molecules are oriented towards the membrane interior, consistent with experimental observations. [23] In contrast, within Region 1 defined by $z^{\prime}<-0.25$ nm, water molecules tend to orient upward, and this tendency becomes more pronounced with increasing $x_{\mathrm{Chol}}$.
To highlight how their dipole alignment varies with $x_{\mathrm{Chol}}$, Fig. 6 plots the average orientation $\langle\cos\theta\rangle$ of water molecules as a function $z^{\prime}$ at 303 K. The results at 323 K are illustrated in Fig. S3 of the supplementary material. As mentioned above, in Region 2, water molecules tend to orient their H-atoms toward the lipid membrane; however, the tendency weakens for $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$. This reduction is likely due to the increase in the lateral spacing between lipid headgroups, as discussed in Sec. III.2. Interestingly, at $x_{\mathrm{Chol}}=0.1$, DPPC exhibited a slight increase in water molecule orientation toward the membrane interior, attributed to the condensation effect of Chol, whereas PSM showed almost no change compared to the $x_{\mathrm{Chol}}=0$ case. These results are consistent with the observed $\rho(z^{\prime})$ dependence on $x_{\mathrm{Chol}}$ [see Fig. 4(a) and (b)]. In contrast, within the interior region (Region 1) of the membrane, water molecules tend to orient upward toward the membrane, particularly at $x_{\mathrm{Chol}}=0.5$. This behavior can be attributed to the increased probability of H-bonding, with Chol’s hydroxyl group [denoted $\mathrm{O}^{\mathrm{Chol}}$ in Fig. 1(c)] acting as the acceptor and water molecules as the donors.
@@ III.4 Transition dynamics among three regions
To elucidate the dynamics of water molecules, we analyzed their transition behavior across the three defined regions: interior (Region 1), interface (Region 2), and bulk (Region 3), as shown in Fig. 4(a) and (b). As examined in Ref. 55, we calculated $C_{i,j}(t)$ $(i\neq j)$, the conditional probability that a water molecule initially located in Region $i$ at time 0 reaches Region $j$ by time $t$, having crossed exactly one regional boundary during this interval. $C_{i,i}(t)$ represents the probability that the water molecule remains in Region $i$ without entering the other regions during the time interval between 0 and $t$. Accordingly, the summation over all possible $j$ regions satisfies $\sum_{j}C_{i,j}(t)=1$, ensuring conservation of the total number of water molecules. Note that as $t$ increases, $C_{i,i}(t)$ decreases while $C_{i,j}(t)$ increases.
@@
Figure 7(a) and (b) shows the results of $C_{1,1}(t)$ and $C_{1,2}(t)$ for DPPC and PSM at 303 K, respectively. The corresponding results at 323 K are shown in Fig. S4 of the supplementary material. When $x_{\mathrm{Chol}}$ increases from 0 to 0.1, the water dynamics slowed down, indicating that the time required to transition into Region 2 became longer. This result is consistent with the finding reported in Ref. 55. In contrast, at $x_{\mathrm{Chol}}=0.5$, the transition from Region 1 to Region 2 was accelerated for both DPPC and PSM. As shown in Fig. 4(a) and (b), the number density of water molecules near the boundary between the two regions (at $z^{\prime}=-0.25$ nm) is larger at $x_{\mathrm{Chol}}=0.5$ than at other $x_{\mathrm{Chol}}$ values. This local increase in water density likely facilitates the dynamics of water molecules along the membrane normal. Comparison between DPPC and PSM reveals that the acceleration effect of Chol on water dynamics is particularly pronounced in PSM, as evidenced by $C_{1,1}(t)$ and $C_{1,2}(t)$. Notably, while water dynamics in the Chol-free system, PSM are slower than those in DPPC, the absence of differences in area per lipid and $\rho(z^{\prime})$ between the two systems at $x_{\mathrm{Chol}}=0.5$, as shown in Fig. 3 and Fig. 4, indicates that the local water dynamics become comparable at high Chol concentrations.
@@
Next, we present the results of $C_{2,1}(t)$, $C_{2,2}(t)$, and $C_{2,3}(t)$ for DPPC and PSM at 303 K in Fig. 8(a) and (b), respectively. The corresponding results at 323 K are shown in Fig. S5 of the supplementary material. As time $t$ increases, the value of $C_{2,1}(t)$ saturated in the order of $x_{\mathrm{Chol}}=0.1$, 0, and 0.5 for both DPPC and PSM. Specifically, at $x_{\mathrm{Chol}}=0.5$, $C_{2,1}(t)$ converges to approximately 0.21, which is about twice the value observed at $x_{\mathrm{Chol}}=0$ This result corresponds to $C_{1,2}(t)$ shown in Fig. 7, indicating that the exchange between the membrane interior region (Region 1) and the interface region (Region 2) becomes faster with increasing $x_{\mathrm{Chol}}$. Conversely, the saturated value of $C_{2,3}(t)$ decreased in the order of $x_{\mathrm{Chol}}=0.1$, 0, and 0.5. In particular, at $x_{\mathrm{Chol}}=0.5$, water molecules in Region 2 are more likely to transition into Region 1, thereby relatively reducing the proportion that moves into the bulk region (Region 3). According to $C_{2,1}(t)+C_{2,2}(t)+C_{2,3}(t)=1$, the proportion remaining in Region 2, represented by $C_{2,2}(t)$, decayed slightly more slowly as $x_{\mathrm{Chol}}$ increased.
@@
These results of $C_{i,j}(t)$ indicate that the hydration state in the interface region (Region 2) is significantly influenced by Chol concentration. In particular, the findings suggest that water molecules in Region 2 become more stable at $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$ and 0.1.
@@ III.5 H-bond lifetime between water molecules and functional groups
Figure 9: Ratio of H-bond lifetimes, $\tau_{\mathrm{HB}}$, at $x_{\mathrm{Chol}}=0.1$ and 0.50.50.5 to those at $x_{\mathrm{Chol}}=0$, denoted as $\tau_{\mathrm{HB(Chol)}}/\tau_{\mathrm{HB(pure)}}$, for (a) DPPC and (b) PSM.
Finally, to elucidate the interfacial dynamics at a more microscopic level, we analyzed the lifetime of H-bonds formed between water molecules as well as between water molecules and functional groups of DPPC and PSM. The identification of H-bond follows the same criteria described in Sec. III.2. The acceptor oxygen of water molecule is denoted as $\mathrm{O}^{\mathrm{w}}$. Note that the donor oxygen atoms are limited to ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{D5}}$, $\mathrm{O}^{\mathrm{D7}}$) for DPPC and ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{S4}}$, $\mathrm{O}^{\mathrm{S5}}$) for PSM, respectively. This selection is based on the significantly lower number of H-bonds involving other oxygen atoms such as ($\mathrm{O}^{\mathrm{D4}}$, $\mathrm{O}^{\mathrm{D6}}$) for DPPC and $\mathrm{O}^{\mathrm{2}}$ for PSM, as confirmed from the joint probability distribution of $r_{\mathrm{OO}}$ and $\beta$ reported in Ref. 55.
@@
[…]
@@
[…]
@@
The values of $\tau_{\mathrm{HB}}$ at $x_{\mathrm{Chol}}=0.1$ and 0.5, normalized by the corresponding value at $x_{\mathrm{Chol}}=0$, are plotted as a function of the acceptor oxygen type in Fig. 9. Figure  9 shows that the H-bond lifetime $\tau_{\mathrm{HB}}$ increases at $x_{\mathrm{Chol}}=0.1$, with a particularly pronounced effect in DPPC at 303 K. This is consistent with our previous findings. [55] In contrast, at $x_{\mathrm{Chol}}=0.5$, $\tau_{\mathrm{HB}}$ decreases significantly for both DPPC and PSM. For DPPC, $\tau_{\mathrm{HB}}$ is reduced to approximately 50% at 303 K and 70% at 323 K compared to Chol-free system. Furthermore, the dependence on the type of acceptor oxygen becomes relatively weak. A similar tendency is observed for PSM, with a more pronounced reduction in $\tau_{\mathrm{HB}}$ at $x_{\mathrm{Chol}}=0.5$, which exhibits comparable values between 303 K and 323 K.
This acceleration effect of Chol on $\tau_{\mathrm{HB}}$ can be attributed to the increased water content in the membrane interface (Region 2) at $x_{\mathrm{Chol}}=0.5$, as discussed previously [see Fig. 4(a) and (b)]. In addition, the more substantial influence of Chol in PSM is likely due to its intrinsically tight membrane packing in the absence of Chol, compared to DPPC, as noted in Sec. III.1. These findings highlight that Chol concentration modulates the dynamics of H-bonds between water molecules and lipid molecules at the membrane interface in a non-monotonic manner, irrespective of the lipid species.
@@ IV Conclusion
Correspondingly, the dynamics of interfacial water characterized by the transition behavior between the three regions and the H-bond lifetime showed the decrease at $x_{\mathrm{Chol}}=0.1$, followed by an acceleration at $x_{\mathrm{Chol}}=0.5$, consistent with previous MD simulations [49, 53] and spectroscopic experiments [13, 22, 23]. In summary, the non-monotonic behavior in dynamics of water molecules near lipid membrane interface arises from the interplay between membrane packing and increased hydration induced by high Chol concentrations.
@@
The supplementary material presents $\rho(z^{\prime})/\rho_{\mathrm{bulk}}$ and the number of H-bonds, and $N_{\mathrm{HB}}/N_{\mathrm{NN}}$ as a function of $z^{\prime}$ at 323 K (Fig. S1); $\rho(z^{\prime},\cos\theta)$ normalized by $\rho(z^{\prime})$ at 323K (Fig. S2); $\langle\cos\theta\rangle$ as a function of $z^{\prime}$ at 323 K (Fig. S3); $C_{1,j}(t)$ at 323 K (Fig. S4); and $C_{2,j}(t)$ at 323 K (Fig. S5).

### v2

@@ III.3 Water orientation near membrane interface
We further analyze how the orientation of water molecules at the membrane interface depends on Chol concentration. Let us introduce water dipole vector $\bm{\mu}$ and the unit vector direced from membrane interface to water phase $\bm{e}_{z}$. We also define the azimuth angle $\theta$ between $\bm{e}_{z}$ and $\bm{\mu}$ given as $\cos\theta=\bm{e}_{z}\cdot\bm{\mu}$. Note that $\cos\theta>0$ and $\cos\theta\leq 0$ indicate that a water molecule is oriented toward the water phase and toward the membrane center, respectively. Previous MD simulations have demonstrated that water orientation varies with the nature of the interacting functional groups. [84, 85, 86, 36, 87, 88, 89] While these results typically represent the $\cos\theta$ dependence on distance $z$ from the membrane center, it should be noted that our analysis uses $z^{\prime}$, a coordinate defined relative to the nearest phosphorus atom. This approach accounts for local fluctuations of the membrane interface and, in particular, enables a more accurate characterization of water orientation near the lipid headgroup region.
@@
Figure 5 demonstrates that in Region 3 (bulk region) defined by $z^{\prime}>0.85$ nm, $\rho(z^{\prime},\cos\theta)$ is approximately 0.5 regardless of $\cos\theta$, indicating an isotropic orientation of water molecules. Furthermore, it is demonstrated that both DPPC and PSM exhibit higher values at $1<\cos\theta<0.5$ in Region 2 defined by $0.25~\mathrm{nm}<z^{\prime}<0.85~\mathrm{nm}$. This Region 2 likely represents water molecules located near the phosphate groups of the lipid headgroups, as discussed in Sec III.2. The predominantly negative values of $\cos\theta$ suggest that the dipole vectors of these water molecules are oriented toward the membrane center, consistent with experimental observations. [24] In contrast, within Region 1 defined by $z^{\prime}<-0.25$ nm, water molecules tend to orient toward the water phase, and this tendency becomes more pronounced with increasing $x_{\mathrm{Chol}}$.
To highlight how their dipole alignment varies with $x_{\mathrm{Chol}}$, Fig. 6 plots the average orientation $\langle\cos\theta\rangle$ of water molecules as a function $z^{\prime}$ at 303 K. The results at 323 K are illustrated in Fig. S7 of the supplementary material. As mentioned above, in Region 2, water molecules tend to orient their H-atoms along the membrane normal toward the membrane center, as illustrated in Fig. 5. However, this orientational tendency becomes weaker at $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$. This reduction is likely due to the increase in the lateral spacing between lipid headgroups, as discussed in Sec. III.2. Interestingly, at $x_{\mathrm{Chol}}=0.1$, DPPC exhibited a slight increase in water molecule orientation toward the membrane center, attributed to the condensation effect of Chol, whereas PSM showed almost no change compared to the $x_{\mathrm{Chol}}=0$ case. These results are consistent with the observed $\rho(z^{\prime})$ dependence on $x_{\mathrm{Chol}}$ [see Fig. 4(a) and (b)]. In contrast, within the interior region (Region 1) of the membrane, water molecules tend to orient toward the water phase, particularly at $x_{\mathrm{Chol}}=0.5$. This behavior can be attributed to the increased probability of H-bonding, with Chol’s hydroxyl group [denoted $\mathrm{O}^{\mathrm{Chol}}$ in Fig. 1(c)] acting as the acceptor and water molecules as the donors.
@@ III.4 Transition dynamics among three regions
To elucidate the dynamics of water molecules, we analyzed their transition behavior across the three defined regions: interior (Region 1), interface (Region 2), and bulk (Region 3), as shown in Fig. 4(a) and (b). As examined in Ref. 56, we calculated $C_{i,j}(t)$ $(i\neq j)$, the conditional probability that a water molecule initially located in Region $i$ at time 0 reaches Region $j$ by time $t$, having crossed exactly one regional boundary during this interval. $C_{i,i}(t)$ represents the probability that the water molecule remains in Region $i$ without entering the other regions during the time interval between 0 and $t$. Accordingly, the summation over all possible $j$ regions satisfies $\sum_{j}C_{i,j}(t)=1$, ensuring conservation of the total number of water molecules. Note that as $t$ increases, $C_{i,i}(t)$ decreases while $C_{i,j}(t)$ increases.
@@
Figure 7(a) and (b) shows the results of $C_{1,1}(t)$ and $C_{1,2}(t)$ for DPPC and PSM at 303 K, respectively. The corresponding results at 323 K are shown in Fig. S8 of the supplementary material. When $x_{\mathrm{Chol}}$ increases from 0 to 0.1, the water dynamics slowed down, indicating that the time required to transition into Region 2 became longer. This result is consistent with the finding reported in Ref. 56. In contrast, at $x_{\mathrm{Chol}}=0.5$, the transition from Region 1 to Region 2 was accelerated for both DPPC and PSM. As shown in Fig. 4(a) and (b), the number density of water molecules near the boundary between the two regions (at $z^{\prime}=-0.25$ nm) is larger at $x_{\mathrm{Chol}}=0.5$ than at other $x_{\mathrm{Chol}}$ values. This local increase in water density likely facilitates the dynamics of water molecules along the membrane normal. Comparison between DPPC and PSM reveals that the acceleration effect of Chol on water dynamics is particularly pronounced in PSM, as evidenced by $C_{1,1}(t)$ and $C_{1,2}(t)$. Notably, while water dynamics in the Chol-free system, PSM are slower than those in DPPC, the absence of differences in area per lipid and $\rho(z^{\prime})$ between the two systems at $x_{\mathrm{Chol}}=0.5$, as shown in Fig. 3 and Fig. 4, indicates that the local water dynamics become comparable at high Chol concentrations.
@@
Next, we present the results of $C_{2,1}(t)$, $C_{2,2}(t)$, and $C_{2,3}(t)$ for DPPC and PSM at 303 K in Fig. 8(a) and (b), respectively. The corresponding results at 323 K are shown in Fig. S9 of the supplementary material. As time $t$ increases, the value of $C_{2,1}(t)$ saturated in the order of $x_{\mathrm{Chol}}=0.1$, 0, and 0.5 for both DPPC and PSM. Specifically, at $x_{\mathrm{Chol}}=0.5$, $C_{2,1}(t)$ converges to approximately 0.21, which is about twice the value observed at $x_{\mathrm{Chol}}=0.1$. This result corresponds to $C_{1,2}(t)$ shown in Fig. 7, indicating that the exchange between the membrane interior region (Region 1) and the interface region (Region 2) becomes faster with increasing $x_{\mathrm{Chol}}$. Conversely, the saturated value of $C_{2,3}(t)$ decreased in the order of $x_{\mathrm{Chol}}=0.1$, 0, and 0.5. In particular, at $x_{\mathrm{Chol}}=0.5$, water molecules in Region 2 are more likely to transition into Region 1, thereby relatively reducing the proportion that moves into the bulk region (Region 3). According to $C_{2,1}(t)+C_{2,2}(t)+C_{2,3}(t)=1$, the proportion remaining in Region 2, represented by $C_{2,2}(t)$, decayed slightly more slowly as $x_{\mathrm{Chol}}$ increased.
@@
These results of $C_{i,j}(t)$ indicate that the hydration state in the interface region (Region 2) is significantly influenced by Chol concentration. Specifically, the findings suggest that water molecules in Region 2 become more stable at $x_{\mathrm{Chol}}=0.5$ compared to $x_{\mathrm{Chol}}=0$ and 0.1.
@@ III.5 H-bond lifetime between water molecules and functional groups
Figure 9: Ratio of H-bond lifetimes, $\tau_{\mathrm{HB}}$, at $x_{\mathrm{Chol}}=0.1$ and $0.5$ to those at $x_{\mathrm{Chol}}=0$, denoted as $\tau_{\mathrm{HB(Chol)}}/\tau_{\mathrm{HB(pure)}}$, for (a) DPPC and (b) PSM.
Finally, to elucidate the interfacial dynamics at a more microscopic level, we analyzed the lifetime of H-bonds formed between water molecules as well as between water molecules and functional groups of DPPC and PSM. The identification of H-bond follows the same criteria described in Sec. III.2. Note that the acceptor oxygen atoms are limited to ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{2}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{D5}}$, $\mathrm{O}^{\mathrm{D7}}$) for DPPC and ($\mathrm{O}^{\mathrm{1}}$, $\mathrm{O}^{\mathrm{3}}$, $\mathrm{O}^{\mathrm{S4}}$, $\mathrm{O}^{\mathrm{S5}}$) for PSM, respectively. We also note that, in the case of the H-bond between water and the nitrogen atom in PSM, water acts only as an H-bond acceptor (see Fig. S10 of the supplementary material for the 2D PMF using $r_{\mathrm{OO}}$ and $\beta$). This selection is based on the significantly lower number of H-bonds involving other oxygen atoms such as ($\mathrm{O}^{\mathrm{D4}}$, $\mathrm{O}^{\mathrm{D6}}$) for DPPC and $\mathrm{O}^{\mathrm{2}}$ for PSM, as confirmed from the joint probability distribution of $r_{\mathrm{OO}}$ and $\beta$ reported in Ref. 56.
@@
where $\langle\cdots\rangle$ represents the ensemble average over all pairs $(i,j)$ and initial times. [90, 78, 91] A similar H-bond dynamics analysis was previously reported in polymer–water mixtures. [92]
@@
$\tau_{\mathrm{HB}}=\int^{\infty}_{0}P_{\mathrm{HB}}(t)dt=\dfrac{\tau_{\mathrm{KWW}}}{\beta_{\mathrm{KWW}}}\Gamma\left(\dfrac{1}{\beta_{\mathrm{KWW}}}\right),$ (2)
@@
The values of $\tau_{\mathrm{HB}}$ at $x_{\mathrm{Chol}}=0.1$ and 0.5, normalized by the corresponding value at $x_{\mathrm{Chol}}=0$, are plotted as a function of the acceptor oxygen type in Fig. 9. Figure  9 shows that the H-bond lifetime $\tau_{\mathrm{HB}}$ increases at $x_{\mathrm{Chol}}=0.1$, with a particularly pronounced effect in DPPC at 303 K. This is consistent with our previous findings. [56] In contrast, at $x_{\mathrm{Chol}}=0.5$, $\tau_{\mathrm{HB}}$ decreases significantly for both DPPC and PSM. For DPPC, $\tau_{\mathrm{HB}}$ is reduced to approximately 50% at 303 K and 70% at 323 K compared to Chol-free system. Furthermore, the dependence on the type of acceptor oxygen becomes relatively weak. A similar tendency is observed for PSM, with a more pronounced reduction in $\tau_{\mathrm{HB}}$ at $x_{\mathrm{Chol}}=0.5$, which exhibits comparable values between 303 K and 323 K. Elola and Rodriguez reported that $\tau_{\mathrm{HB}}$ is reduced at $x_{\mathrm{Chol}}=0.3$ compared with $x_{\mathrm{Chol}}=0$ for the DPPC systems at 323 K.[50] Thus, the turning point at which the Chol effect on the H-bond dynamics shifts from deceleration to acceleration would lie between $x_{\mathrm{Chol}}=0.1$ and 0.3.
The acceleration effect of Chol on $\tau_{\mathrm{HB}}$ can be attributed to the increased water content in the membrane interface (Region 2) at $x_{\mathrm{Chol}}=0.5$, as discussed previously [see Fig. 4(a) and (b)]. In addition, the more substantial influence of Chol in PSM is likely due to its intrinsically tight membrane packing in the absence of Chol, compared to DPPC, as noted in Sec. III.1. These findings highlight that Chol concentration modulates the dynamics of H-bonds between water molecules and lipid molecules at the membrane interface in a non-monotonic manner, irrespective of the lipid species.
@@ IV Conclusion
Correspondingly, the dynamics of interfacial water characterized by the transition behavior between the three regions. H-bond lifetime increase at $x_{\mathrm{Chol}}=0.5$, consistent with previous MD simulations [50, 54] and spectroscopic experiments [14, 23, 24]. Furthermore, given that the H-bond dynamics is retarded at $x_{\mathrm{Chol}}=0.1$, we can state that the significance of cholesterol varies non-monotonically depending on its concentration. In summary, The non-monotonic behavior in dynamics of water molecules near lipid membrane interface arises from the interplay between membrane packing and increased hydration induced by high Chol concentrations.
@@
[…]